OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year? Answer: A)10.3 years B)11.7 years C)12.5 years D)13.1 years E)13.7 years This is my last question

OpenStudy (anonymous):

\[600=300(1.06)^x\] solve for x

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

\[600=300(1+\frac{.06}{2})^{2x}\] solve for x

OpenStudy (anonymous):

\[600=300(1.03)^{2x}\] \[2=(1.03)^{2x}\] and right away you see that the 300 and 600 were not really part of the problem doubling time just means doubling time

OpenStudy (anonymous):

\[(1.03)^{2x}=2\] \[2x=\frac{\ln(2)}{\ln(1.03)}\] \[x=\frac{\ln(2)}{2\ln(1.03)}\] then a calculator

OpenStudy (anonymous):

i get 11.7